HeI PHOTOELECTRON SPECTROSCOPY AND ELECTRONIC STRUCTURE OF ALKYL-LITHIUM CLUSTERS

\underline{G. TARCZAY}, G. VASS, G. MAGYARFALVI, and L. SZEPES, Department of General and Inorganic Chemistry, Eötvös University, P.O. Box 32, H-1518 Budapest 112, HUNGARY.

HeI photoelectron spectra of some alkyl-lithium tetrameric and hexameric clusters have been recorded. In the low energy region (ca. 6.0-9.0 eV) bands have been assigned to ionization from the Li-C cluster orbitals. Ionization from the triply and doubly degenerate orbitals of the tetramers (T_{d} symmetry) and hexamers (D_{3d} symmetry), respectively, causes Jahn-Teller (JT) distortion of the clusters. Ab initio quantum chemical calculations have been performed to understand the nature and extent of the JT distortions.